The school stage is rectangular in shape. Its diagonal is 13 m the length of the rectangle is 7 m longer than its width. Find the length and the width of the rectangle
Answer by jimdris(2)

X^2 + (X + 7)^2 = 13^2 (Pythagorean theorem)
X^2 + X^2 + 14X + 49 = 169
2X^2 + 14X = 120
X^2 + 7X = 60
X(X+7) = 60
X = 5 WIDTH OF STAGE
X+ 7 = 12 LENGTH OF STAGE